OpenStudy (anonymous):

convert 0.62 to an proper fraction

jimthompson5910 (jim_thompson5910):

0.62 = 62 hundredths 0.62 = 62/100 0.62 = 31/50

OpenStudy (anonymous):

Thank

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

4 1/4-2 3/4

jimthompson5910 (jim_thompson5910):